OpenStudy (anonymous):

Find dy/dx for sin(xy) =√x2 + y2.

hartnn (hartnn):

to make life simpler, you can first square both sides, and then take the derivative :)

OpenStudy (watchmath):

\[ \begin{align*} \sin^2(xy)&=x^2+y^2\\ 2\sin(xy)\left(y+xy'\right)&=2x+2yy'\\ \left(2x\sin(xy)-2y\right)y'&=2x-2y\sin(xy)\\ y'&=\frac{2x\sin(xy)-2y}{2x-2y\sin(xy)} \end{align*} \]

OpenStudy (charlotte123):

xD

OpenStudy (charlotte123):

shamil ;3

OpenStudy (charlotte123):

u can't hide from me ԅ(≖‿≖ԅ)

OpenStudy (charlotte123):

*disappears*

OpenStudy (cggurumanjunath):

@watchmath check out your last step.

OpenStudy (watchmath):

yes, thanks :), it should be the reciprocal of that

OpenStudy (cggurumanjunath):

:) @Shawna311

OpenStudy (cggurumanjunath):